OpenStudy (anonymous):

medal a pole 14 m high casts a shadow of 10 m . At the same time what will be the height of a tree which casts a shadow of 7 m

OpenStudy (campbell_st):

use ratios of corresponding sides in similar triangles original/shadow \[\frac{14}{10} = \frac{x}{7}\] so \[x = \frac{14 \times 7}{10}\]

OpenStudy (atlas):

The ratio of the object to its shadow depends on the angle the sun makes with the object. If the angle the sun makes with the object is constant - as it is on a particular point on a day- then this ration will be constant.

OpenStudy (anonymous):

so its directly proportional?

OpenStudy (atlas):

yes

OpenStudy (anonymous):

so we can cross multiply to find out x

OpenStudy (atlas):

yes

OpenStudy (anonymous):

thanks :)

OpenStudy (atlas):

you are welcome